### v1

# Quantum optical nonreciprocity on a magnetic-free photonic chip

Abstract
Breaking reciprocity enables new regimes of light–matter interaction with broad implications for fundamental physics and emerging quantum technologies. Although various approaches have been explored to achieve optical nonreciprocity, realizing it at the single-photon level has remained a major challenge. Here, we demonstrate magnetic-free optical nonreciprocity—including both isolation and circulation—in the quantum regime, enabled by efficient and noiseless all-optical frequency conversion on an integrated photonic chip. Our device preserves the quantum coherence and entanglement of the input photons while delivering exceptional performance parameters, meeting the stringent demands of classical and quantum information systems. This realization of quantum optical nonreciprocity in a scalable photonic platform opens a pathway toward directional quantum communication and noise-resilient quantum networks.
Reciprocity, a cornerstone of Maxwell’s equations, ensures symmetric light transmission and forms the basis of most conventional optical systems. Breaking this fundamental symmetry of electromagnetic reciprocity reveals new physical phenomena Haldane and Raghu (2008); Wang et al. (2009); Fang et al. (2012); Lu et al. (2014) and enables novel functionalities in both classical and quantum regimes. For example, once this symmetry is broken, it becomes possible to engineer nonreciprocal devices such as optical isolators Jalas et al. (2013), which block unwanted backscattering and are critical for stabilizing lasers and amplifiers in communication networks Li (2002). In the quantum regime, control over electromagnetic reciprocity becomes even more significant. Nonreciprocal components protect superconducting qubits from amplifier noise and allow quantum-limited directional amplification in quantum microwave circuits Kamal et al. (2011); Metelmann and Clerk (2015); Kerckhoff et al. (2015); Lecocq et al. (2017); Bernier et al. (2017); Ranzani et al. (2017); Navarathna et al. (2023). They are also increasingly sought after in optical quantum systems Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025); Ren et al. (2022), where they can protect sensitive quantum components from stray light in a network and enable new forms of light–matter interaction for quantum simulation, sensing, and communication Stannigel et al. (2012); Gonzalez-Ballestero et al. (2015); Lodahl et al. (2017); Huang et al. (2018); Barik et al. (2018); Lau and Clerk (2018); Mahmoodian et al. (2020); Ahmadi et al. (2024).
A variety of approaches have been explored to achieve optical nonreciprocity, including magneto-optic effects Bi et al. (2011); Ren et al. (2022) and optical nonlinearities Fan et al. (2012); White et al. (2023). However, magneto-optic materials are intrinsically lossy and difficult to integrate on-chip, while bulk optical nonlinearities are too weak to produce nonreciprocity at the single-photon level. Recently, atomic systems have been shown to be effective for quantum optical nonreciprocity Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025), but they remain constrained by fixed operational wavelengths, narrow bandwidths, and limited scalability. Another widely investigated strategy for inducing optical nonreciprocity is spatiotemporal modulation Yu and Fan (2009); Tzuang et al. (2014); Shen et al. (2016); Fang et al. (2017); Sohn et al. (2021); Tian et al. (2021); Kittlaus et al. (2021); Shah et al. (2023); Cheng et al. (2025); Herrmann et al. (2022); Yu et al. (2023); Sounas and Alù (2017). Unlike passive approaches, spatiotemporal modulation requires external driving fields to break time-reversal symmetry. Whether such actively driven nonreciprocal optical systems can function in the quantum regime—where stringent constraints on loss and noise apply—remains an open question.
Here, we demonstrate the first magnetic-free optical nonreciprocity—encompassing both isolation and circulation—at the single-photon level on an integrated photonic chip, while preserving the quantum coherence and entanglement of the input photons. Our device employs all-optical frequency conversion in nonlinear nanophotonic waveguides with parametric pumps, achieving an added noise well below the single-photon limit. Our platform simultaneously achieves a high extinction ratio of 34 dB, low insertion loss of 0.8 dB, broad bandwidth of 44 GHz, high operational fidelity of 97% (for the circulator), and widely tunable operation wavelengths, significantly outperforming previous demonstrations of quantum optical nonreciprocity based on atomic systems Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025) and magneto-optic effects Ren et al. (2022). These results demonstrate the feasibility of spatiotemporally modulated optical nonreciprocity in the quantum regime and pave the way toward robust, integrated quantum optical systems with broken time-reversal symmetry.
Figure 1: Optical nonreciprocity via noiseless photonic frequency conversion. a. Illustration of optical isolation via photonic frequency conversion in a $\chi^{(2)}$ nonlinear waveguide. b. Representation of $\chi^{(2)}$-mediated optical frequency conversion in the band diagram. c. Optical circulation utilizing the full mode space of a parametrically-pumped $\chi^{(2)}$ waveguide.
Concept
Our quantum optical nonreciprocity is realized on the basis of noiseless parametric frequency conversion Kumar (1990) in a waveguide with $\chi^{(2)}$ nonlinearity (Figs. 1a and b), although the principle can also be extended to systems with $\chi^{(3)}$ nonlinearity. Consider a waveguide supporting three optical modes, $a$, $b$, and $c$, where the $\chi^{(2)}$ nonlinearity facilitates a three-wave mixing process described by the Hamiltonian $\hat{H}=\hbar\chi(\hat{a}^{\dagger}\hat{b}^{\dagger}\hat{c}+\hat{a}\hat{b}\hat{c}^{\dagger})$ with $\hat{a}$ ($\hat{a}^{\dagger}$) denoting the annihilation (creation) operator for mode $a$, and similarly for modes $b$ and $c$. In the forward direction, when the three modes satisfy the frequency- and phase-matching conditions $\omega_{a}+\omega_{b}=\omega_{c}$ and $k_{a}+k_{b}=k_{c}$, where $\omega$ and $k$ denote the angular frequency and wavevector, respectively, and mode $a$ is driven by a classical pump field with amplitude $\alpha$, the interaction between modes $b$ and $c$ becomes linearized, i.e., $\hat{H}=\hbar\chi\alpha(\hat{b}^{\dagger}\hat{c}+\hat{b}\hat{c}^{\dagger})$, which enables noiseless frequency conversion between the two modes, i.e., sum-frequency generation (SFG) ($b\rightarrow c$) and difference-frequency generation ($c\rightarrow b$).
Figure 2: InGaP $\chi^{(2)}$ waveguide with tunable phase matching. a. InGaP photonic chip. b. 3D illustration of the phase-matching tunable waveguide device (not to scale). c. Optical microscope image of an InGaP nanophotonic waveguide integrated with a nanoheater array. d. Scanning electron microscope image of a portion of the device and simulated field profile of the waveguide modes. e. Initial and optimized normalized SFG spectra. Inset shows the voltages of nanoheaters used for optimizing the spectrum.
When the frequency conversion is efficient, the input photons in mode b(c)b\ (c) can be completely transferred to mode c(b)c\ (b). Thus, filtering out the pump and light in mode c(b)c\ (b) effectively blocks the forward transmission of photons in mode b(c)b\ (c). In the backward direction, the same pump configuration no longer satisfies the phase-matching condition, and frequency conversion between modes $b$ and $c$ is suppressed. Consequently, the light input in mode b(c)b\ (c) is transmitted, resulting in nonreciprocal transmission. We note that the definition of forward and backward directions here are opposite to that usually used for an isolator. Notably, the same isolator works simultaneously in widely separated optical frequency bands and inherently integrates with quantum frequency conversion, which is appealing for applications in hybrid quantum networks involving photons of disparate wavelengths. When the full mode space is considered, the scattering matrix of the device, defined by $(b_{\mathrm{L},\mathrm{out}},b_{\mathrm{R},\mathrm{out}},c_{\mathrm{L},\mathrm{out}},c_{\mathrm{R},\mathrm{out}})^{T}=\mathcal{S}(b_{\mathrm{L},\mathrm{in}},b_{\mathrm{R},\mathrm{in}},c_{\mathrm{L},\mathrm{in}},c_{\mathrm{R},\mathrm{in}})^{T}$, is thus asymmetric—a key feature of optical nonreciprocity Jalas et al. (2013):
$\mathcal{S}=\begin{bmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\end{bmatrix}$. (1)
This scattering matrix indicates that the $\chi^{(2)}$ waveguide can also realize a four-port optical circulator when the full mode space is utilized (Fig. 1c), with $(b_{\mathrm{L}},b_{\mathrm{R}},c_{\mathrm{L}},c_{\mathrm{R}})$ corresponding to ports 141-4, respectively. In addition, the directionality of the isolation and circulation can be reversed by switching the propagation direction of the pump.
This form of optical nonreciprocity, which is based on $\chi^{(2)}$- or $\chi^{(3)}$-mediated parametric frequency conversion, is distinct from passive nonlinear optical approaches Fan et al. (2012); White et al. (2023). Because the interaction is linearized by a classical pump, it circumvents the dynamic reciprocity constraint that limits passive nonlinear isolators Shi et al. (2015). Moreover, unlike nonreciprocity via parametric down-conversion Abdelsalam et al. (2020)—which relies on pumping the highest-frequency mode and thus suffers from spontaneous emission noise—our approach leverages noiseless conversion while preserving the original signal frequency in the transmission direction when operating as an isolator. These are critical for the realization of quantum optical isolation, as they avoid the need for single-photon nonlinearities and allow seamless downstream information processing.
Integrated nonlinear photonics platform
For a waveguide of length $L$, parametric frequency conversion causes Rabi oscillation between the two modes governed by the pump power. The isolation ratio is given by (Supplementary Information (SI))
$\mathcal{I}=\frac{\Delta k^{2}}{4g^{2}+\Delta k^{2}}+\frac{4g^{2}}{4g^{2}+\Delta k^{2}}\cos^{2}\sqrt{g^{2}+\frac{\Delta k^{2}}{4}}L,$ (2)
where $g=\sqrt{\frac{\omega_{b}}{\omega_{c}}\eta_{\mathrm{SFG}}P_{a}}$, $\eta_{\mathrm{SFG}}$ is the SFG efficiency in the weak pump limit, and $\Delta k=k_{c}-k_{a}-k_{b}$ is the wavelength-dependent phase mismatch between the optical modes. Maximum isolation occurs at half of the oscillation period for specific pump powers, where the energy transfer between the modes is optimized, with the maximum isolation ratio given by
$\mathcal{I}_{\mathrm{max}}=\left((2n+1)\frac{\pi}{L}\right)^{2}/\Delta k^{2},$ (3)
for integer $n\geq 0$. The maximum isolation ratio is fundamentally constrained by the waveguide’s phase mismatch and is realized when $n+\frac{1}{2}$ oscillations occur between modes $b$ and $c$. Consequently, the essential requirements for achieving effective isolation are minimal phase mismatch and sufficient optical nonlinearity. The latter condition is critical for reducing the required pump power and thereby suppressing pump-induced parasitic noise.
We utilize the InGaP integrated photonics platform (Figs. 2a-b) with substantial $\chi^{(2)}$ nonlinearity ($\chi^{(2)}=220$ pm/V) and low optical losses Zhao and Fang (2022); Akin et al. (2024a, b) to demonstrate optical nonreciprocity (see the SI for device fabrication). The InGaP nanophotonic waveguide is designed to be modal-phase-matched between the 1550-nm-band fundamental transverse electric mode (for $a$ and $b$) and the 780-nm-band fundamental transverse magnetic mode (for $c$). For a phase-matched InGaP nanophotonic waveguide along the (110) direction, the simulated SFG efficiency at the weak pump limit is $\eta_{\mathrm{SFG}}\equiv\frac{P_{c}}{P_{a}P_{b}L^{2}}=520,000\%$/W/cm2. However, because of the nonuniform thickness of the InGaP thin film, phase mismatch accumulates along the waveguide, limiting the frequency conversion efficiency and thus the isolation ratio. To overcome this challenge, we develop an in situ programmable phase-matching tuning method, with nanoheater arrays fabricated adjacent to the waveguide (Figs. 2c-d). The nanoheaters can precisely tune the temperature of each segment of the waveguide and counteract the effect of thickness variations on the phase-matching condition (SI). Fig. 2e shows the initial and optimized SFG spectra of a 6-mm-long waveguide integrated with an array of 15 nanoheaters. The optimized SFG spectrum resembles a sinc2 form, indicating near-perfect phase matching. By minimizing the phase mismatch along the entire waveguide in this way, a high isolation ratio can be achieved.
Figure 3: Isolator performance. a. Backward transmission of the 6-$m$ isolator for pump on and off. b. Measured and theoretical isolation versus pump power. c. Peak isolation ratio versus pump power. d. Tunable isolation by changing the pump wavelength. Two isolation spectra are shown for pump wavelength of 1562 nm and 1543 nm, respectively. The 3-dB and 10-dB isolation bandwidths are indicated. Measured isolation wavelength range is 1538 nm-1570 nm due to the available laser wavelengths while the simulated isolation range is >1000>1000 nm for the same device.
Isolator performance
[…]
Figure 4: Quantum optical isolator and circulator. a. Forward and backward on-chip noise flux in the 1550-nm band at the wavelength separated from the pump by 20 nm (2.5-2.5 THz) versus the pump power of the 6-$m$ waveguide. The line is a linear fitting. b. Normalized correlation function $g^{(2)}(\tau)$ of non-degenerate correlated photon pairs in the forward direction for pump on and off. EPS: entangled-photon source. Line is fitting. c. Two-photon interference fringe of degenerate correlated photon pairs after transmission through the optical isolator in the backward direction. Time-resolved coincidence counts are collected in 250-ps time-bins and for 180 s. The visibility of the fringe is $91.7\%$. d. Measured transmission matrix in the counterclockwise direction. The operational fidelity is $\mathcal{F}=0.962(4)$. e. Measured transmission matrix in the clockwise direction. The operational fidelity is $\mathcal{F}=0.970(3)$. f. Total and noise-subtracted coincidence counts in 2.3 GHz bandwidth in 60 s versus entangled photon pair rate after the signal is transmitted from port 1 to port 4 in the counterclockwise circulator.
The isolation wavelengths, $\lambda_{b}$ and $\lambda_{c}$, of the dual-band isolator can be tuned by the pump wavelength, $\lambda_{p}$, since the isolation is realized via the frequency- and phase-matching conditions of the frequency conversion: $\frac{1}{\lambda_{p}}+\frac{1}{\lambda_{b}}=\frac{1}{\lambda_{c}}$ and $\frac{n_{p}}{\lambda_{p}}+\frac{n_{b}}{\lambda_{b}}=\frac{n_{c}}{\lambda_{c}}$, where $n_{k}$ is the effective index of the respective mode and wavelength. Through numerical simulation, we find wide isolation wavelength ranges of >1000>1000 nm and >30>30 nm exist for modes $b$ and $c$, respectively, for the fabricated waveguide. Owing to the available laser wavelengths in the experiment, to demonstrate the tuning of the isolation wavelength, we set the pump at 1562 nm and reoptimize the phase-matching condition of the waveguide via the nanoheater array. In this case, a maximum isolation of 30.3 dB is achieved at 1546.7 nm with a pump power of 29 mW, as shown in Fig. 3d.
Quantum isolator and circulator
Quantum frequency conversion mediated by the beamsplitter-like Hamiltonian is fundamentally noiseless Kumar (1990). However, parasitic noise due to the pump–matter interaction in the waveguide can still arise. The large optical nonlinearity of InGaP effectively reduce the pump power and suppress parasitic noise. The measured on-chip noise is linear in pump power and reaches approximately $2-3\times 10^{-4}$ counts per second per Hz (cps/Hz) for both backward and forward directions (Fig. 4a) at the signal wavelength, i.e., approximately 20 nm away from the pump, and for the pump power corresponding to the maximum isolation. This linear noise is due to spontaneous Raman scattering of the pump (SI). The noise is well below the single-photon level of approximately 1 cps/Hz, i.e., one photon per time–frequency mode, indicating that our optical isolator can operate in the quantum regime. We note the measured noise is not at the fundamental limit and can be further reduced by separating signal from the pump beyond the Raman peaks of the waveguide material.
To demonstrate the quantum-compatibility of our device, we first verify that signal isolation works for single photons. We use a pair of correlated non-degenerate photons in the telecom band with a bandwidth of 2.3 GHz generated from spontaneous parametric down-conversion (SPDC) and send the signal photon in the forward direction of the isolator. Fig. 4b shows the coincidence counts of the signal and idler photons when the pump is switched on and off. When the pump is off, the measured coincidence peak verifies the strong correlation between signal and idler; when the pump is on, signal photons are block (by conversion to the 780-nm band and filtering), and consequently the coincidence vanishes. We then demonstrate that the time–energy entanglement of the SPDC photons is preserved after transmission through the optical isolator. As illustrated in Fig. 4c, we send a pair of degenerate SPDC photons with a bandwidth of 0.95 nm in the backward direction of the optical isolator and measure the time-resolved two-photon interference Brendel et al. (1991) via an unbalanced glass Mach–Zehnder interferometer (MZI). The MZI has a path delay of 1 ns, which is much longer than the coherence time of the SPDC photons, and its phase difference is controlled through temperature. The measured visibility of the two-photon interference fringe is 91.7% limited by the imperfect interferometer visibility, which exceeds the Clauser–Horne limit of $1/\sqrt{2}\approx 70.7\%$ Clauser and Horne (1974) and thereby confirms the time–energy entanglement of the transmitted photons (see the SI for measurement details).
We then utilize both modes $b$ and $c$ of the $\chi^{(2)}$ waveguide and the quantum frequency conversion between them to achieve optical circulation (see Fig. 1c). In this experiment, we use a 3.2-mm-long phase-matching-optimized waveguide. To characterize the optical circulation, we measure the transmission matrix $T_{ij}\equiv\langle\hat{F}_{i,\mathrm{out}}\rangle/\langle\hat{F}_{j,\mathrm{in}}\rangle$, where $\langle\hat{F}_{i,\mathrm{in(out)}}\rangle$ represents the input (output) photon flux of port $i$. Fig. 4d displays the measured transmission matrix for the four-port circulator with a 31-mW pump power (SI). To quantify the performance of the circulator, we calculate the operational fidelity Scheucher et al. (2016)
$\mathcal{F}=\frac{\mathrm{Tr}[\tilde{T}\cdot T_{\mathrm{id}}^{T}]}{\mathrm{Tr}[T_{\mathrm{id}}\cdot T_{\mathrm{id}}^{T}]},$ (4)
where $\tilde{T}$ is the renormalized transmission matrix, $\tilde{T}_{ij}=T_{ij}/\sum_{k}T_{kj}$, and $T_{\mathrm{id}}$ is the transmission matrix of an ideal circulator. It can be shown that $0\leq\mathcal{F}\leq 1$ and for reciprocal transmission ($\tilde{T}=\tilde{T}^{T}$), the fidelity is bounded by $\mathcal{F}\leq 0.5$ (SI). The fidelity of the measured circulator is $\mathcal{F}=0.962(4)$, while the average insertion loss is $\eta=-10\mathrm{log}\left(\frac{1}{4}\sum_{ij}T_{ij}\right)=2.8$ dB, which is limited by the 780-nm-band waveguide loss. The circulation direction can be reversed by changing the pump direction, and the measured transmission matrix for this case is shown in Fig. 4e with a fidelity of $\mathcal{F}=0.970(3)$. The added noise due to spontaneous Raman scattering of the pump is $7.3\times 10^{-5}$ cps/Hz in the 1550-nm band and $2.5\times 10^{-5}$ cps/Hz in the 780-nm band via up-conversion, substantially lower than the single-photon level, indicating the circulator preserves quantum coherence and entanglement of the input photons. To explicitly demonstrate this, we use a pair of correlated non-degenerate photons in the telecom band with a bandwidth of 2.3 GHz and only send the signal photon in port 1. We measure the correlation between the up-converted signal in port 4 and the idler. The noise-subtracted coincidence counts ($C_{0}$) and total coincidence counts ($C_{t}$) in the 2.3 GHz bandwidth are displayed in Fig. 4f. Based on this, the fidelity of the entangled photons after the circulator is found to be $F=\frac{1}{2}(1+\frac{C_{0}}{C_{t}})=0.968(6)$ (SI).
Table 1: Quantum optical isolator performances
Scheme Isolation Bandwidth Insertion loss Pump power Reference
Quantum frequency conversion 34 dB 35.5 GHz 0.8 dB 27 mW This work
Magneto-optic effect 25 dB 1.6 THz 20 dB / Ref. Ren et al. (2022)
Single Rb atom 13 dB 230 MHz 1.4 dB / Ref. Sayrin et al. (2015)
Rb atomic gas 22.5 dB 200 MHz 1.95 dB 100 mW Ref. Dong et al. (2021)
Rb atomic gas 30.3 dB 175 MHz 0.6 dB 40 mW Ref. Zhang et al. (2023)
Table 2: Quantum optical circulator performances
Scheme Fidelity Average isolation Bandwidth Insertion loss Reference
Quantum frequency conversion 0.97 26 dB 48.8 GHz 2.8 dB This work
Single Rb atom 0.73 7 dB 0.23 GHz 1.4 dB Ref. Scheucher et al. (2016)
Outlook
In conclusion, we demonstrated a new form of optical nonreciprocity based on optical parametric frequency conversion that operates across a wide range of light levels—from individual photons to milliwatt powers (SI)—and spans a broad spectral range, meeting the demands of both classical and quantum information systems. Our integrated quantum optical isolator and circulator provide exceptional performance parameters, significantly outperforming previous demonstrations based on atomic systems and magneto-optic effects (see Tables 1 and 2). Our approach, implemented directly on a scalable platform, enables the construction of robust classical and quantum optical circuits. For example, arranging multiple four-port circulators in one- or two-dimensional arrays can realize programmable networks that guide light in specific directions. The same strategy can be extended to other photonic materials, including those hosting quantum emitters (e.g., quantum-dot-embedded III-V semiconductors), providing a pathway to on-chip quantum light–matter interfaces with broken time–reversal symmetry Gonzalez-Ballestero et al. (2015); Mahmoodian et al. (2020); De Bernardis et al. (2023). With inherent reconfigurability and quantum frequency conversion, this platform opens new opportunities for directional quantum communication and noise-resilient quantum networks Lodahl et al. (2017); Ahmadi et al. (2024); Almanakly et al. (2025).
Acknowledgements
This work is supported by US National Science Foundation under Grant No. 2223192 and QLCI-HQAN (Grant No. 2016136) and U.S. Department of Energy Office of Science National Quantum Information Science Research Centers. S. F. acknowledges the support of a MURI project from the Air Force Office of Scientific Research (Grant No. FA9550-22-1-0339).
Author contributions
K.F. conceived the experiment. J.A., J.H. designed the device. J.A., J.H. fabricated the device. J.H., H.Y., J.A. performed the experiment and analyzed the data. All authors contributed to the writing of the paper.