### v2

# Quantum-coherent optical isolation and circulation using frequency conversion on a chip

Abstract
Breaking optical reciprocity enables new regimes of light–matter interaction with broad implications for fundamental physics and emerging quantum technologies. Although various approaches have been explored to achieve optical nonreciprocity, realizing it at the single-photon level has remained a major challenge. Here, we demonstrate nonmagnetic optical nonreciprocity—including both isolation and circulation—in the quantum regime, enabled by efficient and noiseless all-optical frequency conversion on an integrated III-V photonic chip. Our device preserves the quantum coherence and entanglement of the input photons while delivering exceptional performance parameters, including a high extinction ratio of 34 dB, low insertion loss of 0.8 dB, broad bandwidth of 44 GHz, high operational fidelity of 97%, and widely tunable operation wavelength. This realization of quantum optical nonreciprocity in a scalable photonic platform opens a pathway toward directional quantum communication and noise-resilient quantum networks.
Reciprocity, a cornerstone of Maxwell’s equations, ensures symmetric light transmission and forms the basis of most conventional optical systems. Breaking this fundamental symmetry of electromagnetic reciprocity reveals new physical phenomena Haldane and Raghu (2008); Wang et al. (2009); Fang et al. (2012); Lu et al. (2014) and enables novel functionalities in both classical and quantum regimes. For example, once this symmetry is broken, it becomes possible to engineer nonreciprocal devices—components that allow light to transmit preferentially in one direction—such as optical isolators Jalas et al. (2013), which block unwanted backscattering and are critical for stabilizing lasers and amplifiers in communication networks Li (2002). In the quantum regime, the ability to control electromagnetic reciprocity becomes even more significant, as nonreciprocal components can protect superconducting qubits from amplifier noise and enable quantum-limited directional amplification in quantum microwave circuits Kamal et al. (2011); Metelmann and Clerk (2015); Kerckhoff et al. (2015); Lecocq et al. (2017); Bernier et al. (2017); Ranzani et al. (2017); Navarathna et al. (2023). In optical quantum systems, nonreciprocal devices are increasingly sought after Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025); Ren et al. (2022), where they can protect sensitive quantum components from stray light in a network and enable new forms of light–matter interaction for quantum simulation, sensing, and communication Stannigel et al. (2012); Gonzalez-Ballestero et al. (2015); Lodahl et al. (2017); Huang et al. (2018); Barik et al. (2018); Lau and Clerk (2018); Mahmoodian et al. (2020); Ahmadi et al. (2024).
A variety of approaches have been explored to achieve optical nonreciprocity, including magneto-optic effects Bi et al. (2011); Ren et al. (2022) and passive optical nonlinearities Fan et al. (2012); White et al. (2023). However, magneto-optic materials are intrinsically lossy and difficult to integrate on-chip, while passive optical nonlinearities are too weak to produce nonreciprocity at the single-photon level. Recently, atomic systems have been shown to be effective for quantum optical nonreciprocity Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025), but they remain constrained by fixed operational wavelengths, narrow bandwidths, and limited scalability. Another widely investigated strategy for inducing optical nonreciprocity is spatiotemporal modulation Yu and Fan (2009); Tzuang et al. (2014); Shen et al. (2016); Fang et al. (2017); Kittlaus et al. (2018); Sohn et al. (2021); Tian et al. (2021); Kittlaus et al. (2021); Cheng et al. (2025); Shah et al. (2023); Herrmann et al. (2022); Yu et al. (2023); Sounas and Alù (2017). Unlike passive approaches, spatiotemporal modulation requires external driving fields to break time-reversal symmetry. Whether such actively driven nonreciprocal optical systems can function in the quantum regime—where stringent constraints on loss and noise apply—remains an open question.
Here, we demonstrate the first nonmagnetic optical nonreciprocity—encompassing both isolation and circulation—at the single-photon level on an integrated photonic chip, while preserving the quantum coherence and entanglement of the input photons. Our device employs optical frequency conversion in nonlinear nanophotonic waveguides with parametric optical pumps, achieving an added noise far below the single-photon limit. Thanks to the efficient optical frequency conversion, the pump power required for high isolation is significantly lower than that of spatiotemporally modulated isolators using electro-optic or acousto-optic effects Cheng et al. (2025); Tian et al. (2021); Sohn et al. (2021); Kittlaus et al. (2021, 2018); Shah et al. (2023); Yu et al. (2023); Herrmann et al. (2022). Our platform simultaneously achieves a high extinction ratio of 34 dB, low insertion loss of 0.8 dB, broad bandwidth of 44 GHz, high operational fidelity of 97% (for the circulator), and widely tunable operation wavelengths, significantly outperforming previous demonstrations of quantum optical nonreciprocity based on atomic systems Sayrin et al. (2015); Scheucher et al. (2016); Dong et al. (2021); Hu et al. (2021); Zhang et al. (2023); Wang et al. (2025); Zhang et al. (2025) and magneto-optic effects Ren et al. (2022). These results demonstrate the feasibility of spatiotemporally modulated optical nonreciprocity in the quantum regime and advance the path toward robust, integrated quantum optical systems with broken time-reversal symmetry. Our approach based on parametrically driven optical nonlinearity also provides a solution for integrated isolators on III-V semiconductor platforms, which are constrained by weak electro-optic and piezoelectric effects.
Figure 1: Optical nonreciprocity via noiseless optical frequency conversion. a. Illustration of optical isolation via optical frequency conversion in a $\chi^{(2)}$ nonlinear waveguide. b. Representation of $\chi^{(2)}$-mediated optical frequency conversion in the band diagram. Only the forward direction satisfies the frequency- and phase-matching conditions for the frequency conversion. c. Optical circulation utilizing the full mode space of a parametrically-pumped $\chi^{(2)}$ waveguide. d. InGaP photonic chip. e. 3D illustration of the phase-matching tunable waveguide device (not to scale). f. Optical microscope image of an InGaP nanophotonic waveguide integrated with a nanoheater array. g. Scanning electron microscope image of a portion of the device and simulated field profile of the waveguide modes.
Concept
Our quantum optical isolator and circulator are realized on the basis of noiseless parametric frequency conversion Kumar (1990) in a waveguide with $\chi^{(2)}$ nonlinearity (Figs. 1a and b), although the principle can also be extended to systems with $\chi^{(3)}$ nonlinearity. Consider a waveguide supporting three optical modes, $a$, $b$, and $c$, where the $\chi^{(2)}$ nonlinearity facilitates a three-wave mixing process described by the Hamiltonian $\hat{H}=\hbar\chi(\hat{a}^{\dagger}\hat{b}^{\dagger}\hat{c}+\hat{a}\hat{b}\hat{c}^{\dagger})$ with $\hat{a}$ ($\hat{a}^{\dagger}$) denoting the annihilation (creation) operator for mode $a$, and similarly for modes $b$ and $c$. In the forward direction, when the three modes satisfy the frequency- and phase-matching conditions $\omega_{a}+\omega_{b}=\omega_{c}$ and $k_{a}+k_{b}=k_{c}$, where $\omega$ and $k$ denote the angular frequency and wavevector, respectively, and mode $a$ is driven by a classical pump field with amplitude $\alpha$, the interaction between modes $b$ and $c$ becomes linearized, i.e., $\hat{H}=\hbar\chi\alpha(\hat{b}^{\dagger}\hat{c}+\hat{b}\hat{c}^{\dagger})$, which enables noiseless frequency conversion between the two modes, i.e., sum-frequency generation (SFG) ($b\rightarrow c$) and difference-frequency generation ($c\rightarrow b$).
Figure 2: Isolator performance. a. Backward and forward transmissions of the 6-$m$ isolator for pump on and off with pump power of 21.3 mW. b. Measured and theoretical isolation versus pump power. c. Peak isolation ratio versus pump power. d. Tunable isolation by changing the pump wavelength. Two isolation spectra are shown for pump wavelength of 1562 nm and 1543 nm, respectively. The 3-dB and 10-dB isolation bandwidths are indicated. Measured isolation wavelength range is 1538 nm-1570 nm due to the available laser wavelengths while the simulated isolation range is >1000>1000 nm for the same device.
When the frequency conversion is efficient, the input photons in mode b(c)b\ (c) can be completely transferred to mode c(b)c\ (b). Thus, filtering out the pump and light in mode c(b)c\ (b) effectively blocks the forward transmission of photons in mode b(c)b\ (c). In the backward direction, the same pump configuration no longer satisfies the phase-matching condition, and frequency conversion between modes $b$ and $c$ is suppressed. Consequently, the light input in mode b(c)b\ (c) is transmitted, resulting in nonreciprocal transmission. It should be noted that in this context, the definitions of forward and backward directions are opposite to those typically used for an isolator. Importantly, this isolator operates simultaneously in widely separated optical frequency bands and inherently integrates quantum frequency conversion, making it particularly attractive for hybrid quantum networks involving photons of disparate wavelengths. When the full mode space is considered, the scattering matrix of the device, defined by $(b_{\mathrm{L},\mathrm{out}},b_{\mathrm{R},\mathrm{out}},c_{\mathrm{L},\mathrm{out}},c_{\mathrm{R},\mathrm{out}})^{T}=\mathcal{S}(b_{\mathrm{L},\mathrm{in}},b_{\mathrm{R},\mathrm{in}},c_{\mathrm{L},\mathrm{in}},c_{\mathrm{R},\mathrm{in}})^{T}$, becomes asymmetric—a key feature of optical nonreciprocity Jalas et al. (2013):
$\mathcal{S}=\begin{bmatrix}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 1&0&0&0\end{bmatrix}$. (1)
This scattering matrix indicates that the $\chi^{(2)}$ waveguide can also function as a four-port optical circulator when all modes are utilized (Fig. 1c), with $(b_{\mathrm{L}},b_{\mathrm{R}},c_{\mathrm{L}},c_{\mathrm{R}})$ corresponding to ports 141-4, respectively. Furthermore, the directionality of both isolation and circulation can be reversed by switching the propagation direction of the pump.
This form of optical nonreciprocity, which is based on $\chi^{(2)}$- or $\chi^{(3)}$-mediated parametric frequency conversion, is distinct from passive nonlinear optical approaches Fan et al. (2012); White et al. (2023). Because the interaction is linearized by a classical pump, it circumvents the dynamic reciprocity constraint that limits passive nonlinear isolators Shi et al. (2015). Moreover, unlike nonreciprocity via parametric down-conversion Abdelsalam et al. (2020)—which relies on pumping the highest-frequency mode and thus suffers from spontaneous emission noise—our approach leverages noiseless conversion while preserving the original signal frequency in the transmission direction when operating as an isolator. These are critical for the realization of quantum optical isolation, as they avoid the need for single-photon nonlinearities and allow seamless downstream information processing.
Integrated nonlinear photonics platform
For a waveguide of length $L$, parametric frequency conversion causes Rabi oscillation between the two modes governed by the pump power. The isolation ratio is given by (Supplementary Information (SI))
$\mathcal{I}=\frac{\Delta k^{2}}{4g^{2}+\Delta k^{2}}+\frac{4g^{2}}{4g^{2}+\Delta k^{2}}\cos^{2}\sqrt{g^{2}+\frac{\Delta k^{2}}{4}}L,$ (2)
where $g=\sqrt{\frac{\omega_{b}}{\omega_{c}}\eta_{\mathrm{SFG}}P_{a}}$, $\eta_{\mathrm{SFG}}$ is the SFG efficiency in the weak pump limit, and $\Delta k=k_{c}-k_{a}-k_{b}$ is the wavelength-dependent phase mismatch between the optical modes. Maximum isolation occurs at half of the oscillation period for specific pump powers, where the energy transfer between the modes is optimized, with the maximum isolation ratio given by
$\mathcal{I}_{\mathrm{max}}=\left((2n+1)\frac{\pi}{L}\right)^{2}/\Delta k^{2},$ (3)
for integer $n\geq 0$. The maximum isolation ratio is fundamentally constrained by the waveguide’s phase mismatch and is realized when $n+\frac{1}{2}$ oscillations occur between modes $b$ and $c$.
Figure 3: Quantum optical isolator and circulator. a. Forward and backward on-chip noise flux in the 1550-nm band at the wavelength separated from the pump by 20 nm (2.5-2.5 THz) versus the pump power of the 6-$m$ waveguide. The line is a linear fitting. b. Normalized correlation function $g^{(2)}(\tau)$ of non-degenerate correlated photon pairs in the forward direction for pump on and off. EPS: entangled-photon source. Line is fitting. c. Two-photon interference fringe of degenerate correlated photon pairs after transmission through the optical isolator in the backward direction. Time-resolved coincidence counts are collected in 250-ps time-bins and for 180 s. The visibility of the fringe is $91.7\%$. d. Measured transmission matrix in the counterclockwise direction. The operational fidelity is $\mathcal{F}=0.962(4)$. e. Measured transmission matrix in the clockwise direction. The operational fidelity is $\mathcal{F}=0.970(3)$. f. Total and noise-subtracted coincidence counts in 2.3 GHz bandwidth in 60 s versus entangled photon pair rate after the signal is transmitted from port 1 to port 4 in the counterclockwise circulator.
We demonstrate the optical nonreciprocity using the III-V InGaP integrated photonics platform (Figs. 1d-e) with a substantial $\chi^{(2)}$ nonlinearity ($\chi^{(2)}=220$ pm/V) and low optical losses Zhao and Fang (2022); Akin et al. (2024a, b) (see the SI for device fabrication). The InGaP nanophotonic waveguide is designed to be modal-phase-matched between the 1550-nm-band fundamental transverse electric mode (for $a$ and $b$) and the 780-nm-band fundamental transverse magnetic mode (for $c$). For a phase-matched InGaP nanophotonic waveguide along the (110) direction, the simulated SFG efficiency at the weak pump limit is $\eta_{\mathrm{SFG}}\equiv\frac{P_{c}}{P_{a}P_{b}L^{2}}=520,000\%$/W/cm2. We also implemented in situ phase-matching tuning nanoheater arrays (Figs. 1f-g) to control the temperature of each segment of the waveguide and counteract the effect of thickness variations on the phase-matching condition (SI). By minimizing the phase mismatch along the entire waveguide, a high isolation ratio can be achieved.
Isolator performance
[…]
The isolation wavelengths, $\lambda_{b}$ and $\lambda_{c}$, of the dual-band isolator can be tuned by adjusting the pump wavelength, $\lambda_{p}$, since the isolation is determined by the frequency- and phase-matching conditions of the frequency conversion: $\frac{1}{\lambda_{p}}+\frac{1}{\lambda_{b}}=\frac{1}{\lambda_{c}}$ and $\frac{n_{p}}{\lambda_{p}}+\frac{n_{b}}{\lambda_{b}}=\frac{n_{c}}{\lambda_{c}}$, where $n_{k}$ is the effective index of the respective mode and wavelength. Numerical simulations indicate that the fabricated waveguide supports wide isolation wavelength ranges, with mode $b$ tunable over more than 1000 nm and mode $c$ over more than 30 nm. Due to the available laser wavelengths in our experiment, we demonstrate the tuning of the isolation wavelength by setting the pump at 1562 nm and reoptimizing the phase-matching condition of the waveguide using the nanoheater array. Under these conditions, a maximum isolation of 30.3 dB is achieved at 1546.7 nm with a pump power of 29 mW, as shown in Fig. 2d.
Quantum isolator and circulator
Quantum frequency conversion mediated by the beamsplitter-like Hamiltonian is fundamentally noiseless Kumar (1990). However, parasitic noise due to pump–matter interactions in the waveguide can still arise. We measure the on-chip noise to be linear in pump power, reaching approximately $2\times 10^{-4}$ counts per second per Hz (cps/Hz) for both backward and forward directions at the signal wavelength, about 20 nm away from the pump, and at the pump power corresponding to maximum isolation (Fig. 3a). This noise level is well below the single-photon threshold of approximately 1 cps/Hz (one photon per time–frequency mode), confirming that our optical isolator can operate in the quantum regime.
To demonstrate the quantum-compatibility of our device, we first verify that signal isolation works for single photons. We use a pair of correlated non-degenerate photons in the telecom band with a bandwidth of 2.3 GHz generated from spontaneous parametric down-conversion (SPDC) and send the signal photon in the forward direction of the isolator. Fig. 3b shows the coincidence counts of the signal and idler photons when the pump is switched on and off. When the pump is off, the measured coincidence peak verifies the strong correlation between signal and idler; when the pump is on, the signal photons are blocked (by conversion to the 780-nm band and subsequent filtering), and consequently the coincidence vanishes. We then demonstrate that the time–energy entanglement of the SPDC photons is preserved after transmission through the optical isolator. As illustrated in Fig. 3c, we send a pair of degenerate SPDC photons with a bandwidth of 0.95 nm in the backward direction of the optical isolator and measure the time-resolved two-photon interference Brendel et al. (1991) via an unbalanced glass Mach–Zehnder interferometer (MZI). The MZI has a path delay of 1 ns, much longer than the coherence time of the SPDC photons, and its phase difference is controlled through temperature. The measured visibility of the two-photon interference fringe is 91.7% limited by the imperfect interferometer visibility, which exceeds the Clauser–Horne limit of $1/\sqrt{2}\approx 70.7\%$ Clauser and Horne (1974), thereby confirming the time–energy entanglement of the transmitted photons (see the SI for measurement details).
Figure 4: Comparison of integrated nonmagnetic isolators. a. Isolation vs. pump power. b. 3-dB isolation bandwidth vs. insertion loss. Acousto-optic effect: Cheng et al. (2025); Tian et al. (2021); Sohn et al. (2021); Kittlaus et al. (2021, 2018). Electro-optic effect: Shah et al. (2023); Yu et al. (2023); Herrmann et al. (2022). Passive Kerr nonlinearity: White et al. (2023).
We then utilize both modes $b$ and $c$ of the $\chi^{(2)}$ waveguide and the quantum frequency conversion between them to achieve optical circulation (see Fig. 1c). In this experiment, we use a 3.2-mm-long phase-matching-optimized waveguide. To characterize the optical circulation, we measure the transmission matrix $T_{ij}\equiv\langle\hat{F}_{i,\mathrm{out}}\rangle/\langle\hat{F}_{j,\mathrm{in}}\rangle$, where $\langle\hat{F}_{i,\mathrm{in(out)}}\rangle$ represents the input (output) photon flux of port $i$. Fig. 3d displays the measured transmission matrix for the four-port circulator with a 31-mW pump power (SI). To quantify the performance of the circulator, we calculate the operational fidelity Scheucher et al. (2016)
$\mathcal{F}=\frac{\mathrm{Tr}[\tilde{T}\cdot T_{\mathrm{id}}^{T}]}{\mathrm{Tr}[T_{\mathrm{id}}\cdot T_{\mathrm{id}}^{T}]},$ (4)
where $\tilde{T}$ is the renormalized transmission matrix, $\tilde{T}_{ij}=T_{ij}/\sum_{k}T_{kj}$, and $T_{\mathrm{id}}$ is the transmission matrix of an ideal circulator. It can be shown that $0\leq\mathcal{F}\leq 1$ and for reciprocal transmission ($\tilde{T}=\tilde{T}^{T}$), the fidelity is bounded by $\mathcal{F}\leq 0.5$ (SI). The fidelity of the measured circulator is $\mathcal{F}=0.962(4)$, while the average insertion loss is $\eta=-10\mathrm{log}\left(\frac{1}{4}\sum_{ij}T_{ij}\right)=2.8$ dB, which is limited by the 780-nm-band waveguide loss. The circulation direction can be reversed by changing the pump direction, and the measured transmission matrix for this case is shown in Fig. 3e with a fidelity of $\mathcal{F}=0.970(3)$. The added noise is $7.3\times 10^{-5}$ cps/Hz in the 1550-nm band and $2.5\times 10^{-5}$ cps/Hz in the 780-nm band via up-conversion, substantially lower than the single-photon level, indicating the circulator preserves quantum coherence and entanglement of the input photons. To explicitly demonstrate this, we use a pair of correlated non-degenerate photons in the telecom band with a bandwidth of 2.3 GHz and only send the signal photon in port 1. We measure the correlation between the up-converted signal in port 4 and the idler. The noise-subtracted coincidence counts ($C_{0}$) and total coincidence counts ($C_{t}$) in the 2.3 GHz bandwidth are displayed in Fig. 3f. Based on this, the fidelity of the entangled photons after the circulator is found to be $F=\frac{1}{2}(1+\frac{C_{0}}{C_{t}})=0.968(6)$ (SI).
Outlook
In conclusion, we demonstrated a new form of optical nonreciprocity based on optical parametric frequency conversion that operates across a wide range of light levels—from individual photons to milliwatt powers—and spans a broad spectral range. Fig. 4 compares key performance parameters of our isolator with other integrated nonmagnetic isolators demonstrated recently: our isolator achieves a remarkable isolation ratio with a low pump power while offering a unique combination of wide bandwidth and low insertion loss. Since III-V semiconductors have weak electro-optic and piezoelectric effects, our approach based on optical nonlinearity provides a solution for integrated isolators on III-V platforms. Moreover, the same strategy can be extended to other nonlinear photonic materials, such as periodically poled LiNbO3, where extremely wide bandwidth of optical frequency conversion (>1>1 THz)—and thus isolation—can be achieved Wang et al. (2023).
In terms of performance at the single-photon level, our integrated quantum optical isolator and circulator also significantly outperform previous demonstrations based on atomic systems and magneto-optic effects (see Tables 1 and 2). Our approach, implemented directly on a scalable platform, enables the construction of robust classical and quantum optical circuits. For example, arranging multiple four-port circulators in one- or two-dimensional arrays can realize programmable networks that guide light in specific directions. Integration with quantum emitters, such as quantum-dot-embedded III-V semiconductors, would allow the construction of on-chip quantum light–matter interfaces with broken time–reversal symmetry Gonzalez-Ballestero et al. (2015); Mahmoodian et al. (2020). With its inherent reconfigurability and quantum frequency conversion capabilities, this platform offers new opportunities for directional quantum communication and noise-resilient quantum networks Lodahl et al. (2017); Ahmadi et al. (2024); Almanakly et al. (2025).
Table 1: Quantum optical isolator performances
Scheme Isolation Bandwidth Insertion loss Pump power Reference
Quantum frequency conversion 34 dB 35.5 GHz 0.8 dB 27 mW This work
Magneto-optic effect 25 dB 1.6 THz 20 dB / Ref. Ren et al. (2022)
Single Rb atom 13 dB 230 MHz 1.4 dB / Ref. Sayrin et al. (2015)
Rb atomic gas 22.5 dB 200 MHz 1.95 dB 100 mW Ref. Dong et al. (2021)
Rb atomic gas 30.3 dB 175 MHz 0.6 dB 40 mW Ref. Zhang et al. (2023)
Table 2: Quantum optical circulator performances
Scheme Fidelity Average isolation Bandwidth Insertion loss Reference
Quantum frequency conversion 0.97 26 dB 48.8 GHz 2.8 dB This work
Single Rb atom 0.73 7 dB 0.23 GHz 1.4 dB Ref. Scheucher et al. (2016)
Acknowledgements
This work is supported by US National Science Foundation under Grant No. 2223192 and QLCI-HQAN (Grant No. 2016136) and U.S. Department of Energy Office of Science National Quantum Information Science Research Centers. S. F. acknowledges the support of a MURI project from the Air Force Office of Scientific Research (Grant No. FA9550-22-1-0339).
Author contributions
K.F. conceived the experiment. J.A., J.H. designed the device. J.A., J.H. fabricated the device. J.H., H.Y., J.A. performed the experiment and analyzed the data. All authors contributed to the writing of the paper.